OpenStudy (watermelon14):

which series below converges? A.2+3+4.5+... B. 100+80+64+... C. 1-2+4-... D. 8+12+18+...

OpenStudy (watermelon14):

@rishavraj ?

satellite73 (satellite73):

hint, at the very least, the numbers have to be getting smaller

OpenStudy (watermelon14):

i am confused with the concept of converge like i know it is when infinite sequences of numbers added to 1

satellite73 (satellite73):

converge means you can add them and get a finite number

satellite73 (satellite73):

if the numbers get bigger and bigger, there is no way for it to converge

OpenStudy (watermelon14):

yeah what does that mean? ex?

satellite73 (satellite73):

you want an example of a sum that converges?

OpenStudy (watermelon14):

yea

satellite73 (satellite73):

ok how about \[\frac{3}{10}+\frac{3}{100}+\frac{3}{1000}+\frac{3}{10000}+...\]

satellite73 (satellite73):

not only does it converge, but you know what it converges to

OpenStudy (watermelon14):

no i don't

satellite73 (satellite73):

you do if you write it as a decimal

satellite73 (satellite73):

\[\frac{3}{10}=0.3\\ \frac{3}{100}=0.03\\ \frac{3}{1000}=0.003\] and so on

OpenStudy (watermelon14):

oh ok

satellite73 (satellite73):

if you add you get \[0.3333...\] which you probably recognize as \(\frac{1}{3}\)

OpenStudy (watermelon14):

yes i see. if i had  how do i get that pattern